### v1

# LED: LLM Enhanced Open-Vocabulary Object Detection without Human Curated Data Generation

Abstract
Large foundation models trained on large-scale visual-text data can significantly enhance Open Vocabulary Object Detection (OVD) through data generation. However, this may lead to biased synthetic data and overfitting to specific configurations. It can sidestep biases of manually curated data generation by directly leveraging hidden states of Large Language Models (LLMs), which is surprisingly rarely explored. This paper presents a systematic method to enhance visual grounding by utilizing decoder layers of the LLM of a MLLM. We introduce a zero-initialized cross-attention adapter to enable efficient knowledge transfer from LLMs to object detectors, an new approach called L ED (LLM Enhanced Open-Vocabulary Object Detection). We demonstrate that intermediate hidden states from early LLM layers retain strong spatial-semantic correlations that are beneficial to grounding tasks. Experiments show that our adaptation strategy significantly enhances the performance on complex free-form text queries while remaining the same on plain categories. With our adaptation, Qwen2-0.5B with Swin-T as the vision encoder improves GroundingDINO by 2.33% on Omnilabel, at the overhead of 8.7% more GFLOPs. Qwen2-0.5B with a larger vision encoder can further boost the performance by 6.22%. We further validate our design by ablating on varied adapter architectures, sizes of LLMs, and which layers to add adaptation.
1 Introduction
Figure 1: Overview of open vocabulary object detection approaches. (a) Data generation methods prone to distribution shift and overfitting; (b) Our proposed adapter-based framework leverages hidden states from the MLLM to provide adaptation prompts to the detection decoder, enhancing semantic richness and generalization without relying on human curated data.
Open Vocabulary object Detection (OVD) localizes objects described by a free-form text, significantly increasing the size of the label space and enabling broader applications, compared to the traditional closed-set object detection with a fixed label space. Great prior efforts [44, 26, 15, 35, 45] focus on exploring large foundation models trained on image and/or text data to generate high-quality data for OVD, greatly boosting the performance on challenge benchmarks [10, 30, 33]. Their success demonstrates that large foundation models implicitly capture rich semantic knowledge and reasoning capabilities through pretraining, which are beneficial for accurately localizing objects under complex, free-form textual queries.
However, existing methods rely on hand-crafted rules to query large foundation models for synthetic data. For example, VL-PLM [44] and OWLv2 [26] design a delicate procedure with heuristics to use pretrained vision and language models (VLMs) to generate pseudo labels for unseen categories that are not included in the human provided data. DesCo [15] and GenEnhancedNegs [46] assume that detailed or negative descriptions are important training signals for OVD, and thus leverage large language models (LLMs) to create those descriptions. Those methods are well designed for certain aspects (e.g. unseen categories and descriptions) to improve OVD models, which are biased and cannot fully leverage knowledge gained during the pretraining of large foundation models.
In this paper, We present L ED , a new approach to leverage the knowledge of large foundation models for OVD without biased hand-crafted rules. Unlike prior studies that explicitly generate data under some assumptions, we believe the intermediate hidden states of large foundation models embody the essential knowledge and develop a novel framework to embed those hidden states into prevailing object detector. Specifically, as shown in Figure 1(a), existing approaches typically generate new data with specific goals, such as generating negative samples to address hallucination issues [37, 41, 29, 36, 46] or generating more detailed descriptions for spatial and logical relationships [15], which in turn limits the robustness and generalization of the model. In contrast as Figure 1(b), we extract hidden states from the LLM of a multimodal LLM (MLLM), which contains rich semantics and knowledge from large-scale pretraining, and integrate them into the object detector.
We apply L ED to popular open-source object detectors. GroundingDino [22]. Our method significantly improves the performance on prevailing benchmarks, including RefCOCO [39] and OmniLabel benchmark [30]. Besides, we find that hidden states from the first several layers of a MLLM are rich enough to improve the OVD models. Probably, MLLM focuses more on vision tokens in early layers and then on text tokens in later layers, as illustrated in section 3. Thus, vision tokens of deeper layers are not quite updated and embody similar knowledge as earlier ones. We believe our finding and analysis are beneficial for future studies that explore hidden states of large foundation models for downstream tasks.
Our contributions are summarized as. (1) We present a new framework that leverages knowledge from large foundation models. Instead of generating data, we use hidden states to improve OVD models. (2) We find and analyze that hidden states of the first several layers of a MLLM contains sufficient knowledge for detectors, which paves the path for future studies exploring hidden states for downstream tasks. (3) We conducted extensive experiments, and show that L ED can improve OVD models on various benchmarks and possesses significant zero-shot capabilities.
2 Relate Work
Multi-Modal Large Language Models: Recent advances in MLLMs have catalyzed research on visual knowledge integration, where modality alignment plays a pivotal role in vision-language learning [11]. CLIP pioneers this direction through contrastive learning on large-scale image-text pairs [27], while BLIP further improves alignment quality via sophisticated data curation strategies [14, 13]. LLaVA introduces visual instruction tuning using GPT-4-generated multimodal data, significantly advancing the general-purpose multimodal reasoning capabilities of large language-and-vision models[20]. Building upon these foundations, InternVL introduces a progressive alignment paradigm that effectively scales vision-language models, achieving state-of-the-art performance in both vision foundation models and VLLMs [4, 5, 6, 9, 32]. LLaMA Adapter introduces a lightweight adaptation method with a zero-initialized attention mechanism, enabling efficient instruction tuning of LLaMA using only 1.2M learnable parameters and achieving performance comparable to fully fine-tuned models [43]. It is noteworthy that LLaMA Adapter improves the performance of specific language tasks by providing the decoder with continuous feature representations of vocabulary as prompts. L ED aims to leverage the vision understanding of MLLM to directly provide high-level feature representation guidance for downstream tasks.
Open-Set Object Detection: While MLLMs excel at integrating textual and visual modalities, open-world detection remains an unsolved challenge that calls for more flexible category representations. DINO advances closed-set detection through contrastive de-noising in DETR frameworks [42, 1], its reliance on fixed category definitions limits open-world applicability. This motivates open-set detection methods that leverage language generalization: OV-DETR employs CLIP embeddings as cross-modal queries for category-agnostic box decoding [40], and GLIP reformulates detection as visual grounding with phrase-region alignment [16, 8]. GroundingDINO subsequently refines these grounding mechanisms, establishing new benchmarks in zero-shot detection [21].
Vision-Language Grounding: With open-set capabilities taking shape, further progress demands robust vision-language grounding that handles diverse and unconstrained data distributions. Recent advances in leveraging vision-language models (VLMs) for open-vocabulary and open-ended object detection predominantly rely on large-scale synthetic or generated datasets. Techniques such as prompt-driven visual-text adaptation [34, 23] and generative augmentation of negative samples [46] have significantly improved model performance by enriching training data with diversified visual and textual examples. Similarly, generative region-language pretraining frameworks [17] and unified object-centric grounding methods [28] also heavily rely on generated or automatically curated datasets, demonstrating their efficiency and scalability.
However, synthetic data often diverges from real-world data distributions, introducing biases that impair model performance under natural variations and unseen conditions. Furthermore, generative techniques relying heavily on constructed datasets may inadvertently overfit specific dataset characteristics, restricting model adaptability when confronted with open-world scenarios or challenging distributions not adequately represented during training. Although methods like GLIP and DetCLIP [34] have enhanced concept grounding through enriched annotations and dictionaries, the intrinsic reliance on synthetic constructs remains a fundamental challenge.
3 Methodology
Our framework establishes an efficient pipeline for transferring multimodal knowledge from Large Language Models (LLMs) to OVD models via intermediate feature adaptation. As shown in Figure 2, we analyzed token attention changes across different layers of the LLM decoder for Qwen2-0.5B. In the early layers of the LLMs decoder (e.g., layers 0 to 4), the median attention score decreases significantly and fluctuates slightly in the subsequent 20 layers. This indicates that the MLLMs already exhibit strong attention or encoding capabilities for input image tokens in the early layers, while the later layers may focus more on fine-tuning task- or semantics-related text tokens. From a practical perspective, this distribution of attention scores suggests that early layers are critical for extracting vision-based features for downstream tasks such as object detection. Meanwhile, later layers can be selectively utilized or fused depending on specific task requirements.
Figure 2: Median distribution of attention scores across different layers in Qwen2-0.5B decoder.
To validate this concept, we designed an architecture as shown in Figure 3. The architecture comprises four key components: (1) An efficient cross-modal alignment strategy in subsection 3.1; (2) A MLLM for semantic adaptation prompts generation in subsection 3.2; (3) A sufficiently universal end-to-end detector in subsection 3.3; (4) An adapter for feature fusion based on zero-initialized cross-attention in subsection 3.4.
Figure 3: The architecture of L ED . We present the overall architecture, MLLM, detector, zero-initialized cross-attention in Block 1, 2, 3, respectively. In stage 1 and 2, a trained projector is used to align the frozen Vision Encoder with the LLM. In stage 3, a zero-initialized adapter is trained to guide the detector with adaptation prompts, while the projector still participate training to adapt to downstream tasks.
3.1 Cross-modal Alignment
L
ED
aims to provide a comprehensive solution by leveraging the extensive knowledge of LLMs to enhance semantic image understanding. In general, MLLMs for text and images require a vision encoder to generate visual tokens for LLMs. Thus, cross-modal alignment is crucial for ensuring that LLMs can interpret the visual tokens generated by the vision encoder and assign semantic meaning to them. To facilitate effective knowledge transfer while maintaining computational efficiency, we establish a shared vision encoder between the MLLM and the detector through a three-stage alignment process:
Stage 1 - Pretraining: We employ a vision encoder and LLM to pretrain the MLP projection layer to align feature spaces.
Stage 2 - Finetuning: We fine-tune the model on a specific task dataset to align the vision encoder and LLM at a fine-grained level and improve instruction-following capabilities.
Stage 3 - Adapter Training: With the MLLM and detector frozen, we train the grounding decoder adapter (Block 3) and the MLP layer and LLM in the MLLM to integrate semantic cues from its intermediate features, thereby enabling the LLM to better focus on grounding tasks when processing visual tokens.
3.2 Semantic Prompt Generation
As illustrated in Figure 3 Block 1, we employ an MLLM that processes the same input images shared with the detector. The MLLM integrates visual features with system prompts and provides the resulting embeddings as input to the LLM.
We extract intermediate representations from the $\ell_{LM}$-th decoder layer of an $n$-layer language model, referred to as the hidden state embedding $\mathbf{H}$. After obtaining the $\mathbf{H}$, the vision embedding $\mathbf{E_{V}}^{L}$ and text embeddings $\mathbf{E_{T}}$ are derived by truncating the embeddings based on the positions of the image and text tokens:
$\mathbf{H}_{\ell_{LM}}=\mathrm{Decoder}_{\ell_{LM}}(\mathbf{H}_{\ell_{LM}-1}),\ell_{LM}\in\{1,...,n\}$ (1)
$\mathbf{E_{V}}^{L},\mathbf{E_{T}}=\text{Truncate}(\mathbf{H}_{\ell_{LM}})$
3.3 Grounding Detector
An End-to-End Transformer-based Detector architecture, as shown in Figure 3 Block 2. For every (Image, Text) pair, this Block employs a vision encoder and a text encoder to independently extract visual and textual features. These features are then passed into a cross-modality decoder, which interacts with both modalities to refine and update the features. The final output queries generated by the last layer of the decoder are used to predict bounding boxes for objects and extract their corresponding textual phrases [21, 42, 1].
3.4 Zero-initialized Cross-attention
An adapter design (Block 3 in Figure 3) effectively refines the vision embedding $\mathbf{E_{V}}^{L}$ and text embedding $\mathbf{E_{T}}$ from the LLM to generate adaptation prompts with detector features through zero-initialized cross-attention. We apply a convolutional layer to process the MLLM’s vision embeddings $\mathbf{E_{V}}^{L}$ to get adaptation prompts $\mathbf{A_{P}}$.
$\mathbf{A_{P}}=\mathrm{Conv2D}(\mathbf{E_{V}}^{L})\in\mathbb{R}^{B\times L\times d}$ (2)
where $B$ is batch size, $L$ is the length of adaptation prompts, decomposed into the height $h$ and width $w$ of the Vision embedding $\mathbf{E_{V}}^{L}$ when performing convolution processing.
A zero-initialized cross-modal attention mechanism is designed based on a small gating unit. It is worth noting that since the weights of the entire adapter are randomly initialized, directly providing the adaptation prompt to the detector decoder introduces disordered noise, which is detrimental to the training of the network. Given an decoder embedding $\mathbf{E_{D}}^{\ell_{D}}\in\mathbb{R}^{B\times T\times d}$ from $\ell_{D}$th decoder layer and adaptation pr ompts $\mathbf{A_{P}}\in\mathbb{R}^{B\times L\times d}$, our zero-initialized cross-modal attention mechanism operates as follows:
$\mathbf{Q}$ $=\mathbf{E_{D}}\mathbf{W}_{Q}\in\mathbb{R}^{B\times T\times h\times d_{h}}$ (3)
$\mathbf{K}$ $=\mathcal{C}(\mathbf{A_{P}}\cdot\mathbf{W}_{K},\mathbf{E_{D}}\mathbf{W}_{K})\in\mathbb{R}^{B\times(L+T)\times h\times d_{h}}$
$\mathbf{V}$ $=\mathcal{C}(\mathbf{A_{P}}\cdot\mathbf{W}_{V},\mathbf{E_{D}}\mathbf{W}_{V})\in\mathbb{R}^{B\times(L+T)\times h\times d_{h}}$
where $\mathcal{C}(\cdot)$ denotes concatenation along the sequence dimension, $h$ is the number of attention heads, and $d_{h}=d/h$ represents the dimension per head. $\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}$ are the linear projection matrices used in the attention mechanism to map to $Q,K,V$.
$\mathbf{S}=$ $\frac{\tilde{\mathbf{Q}}\tilde{\mathbf{K}}^{\top}}{\sqrt{d_{h}}}+\mathbf{M}\in\mathbb{R}^{B\times h\times T\times(L+T)}$ (4)
$\tilde{\mathbf{S}}=$ $\mathcal{C}(\tanh(\mathbf{g})\odot\text{softmax}(\mathbf{S}_{:,0:L}),$
$\text{softmax}(\mathbf{S_{:,L+1:L+T}})\in\mathbb{R}^{B\times h\times T\times(L+T)}$
where RoPE is defined as $\tilde{\mathbf{Q}},\tilde{\mathbf{K}}=\text{RoPE}(\mathbf{Q},\mathbf{K};\Theta)$ with angular parameter matrix $\Theta$; $\mathbf{S}$ represents the raw attention score matrix, and $\tilde{\mathbf{S}}$ is attention weight matrix obtained by applying softmax; $\mathbf{g}\in\mathbb{R}^{1\times h\times 1\times 1}$ is a learnable gating vector initialized to $\mathbf{0}$. and an activation function $\tanh(\mathbf{\cdot})$ is adopted to regulate the scale of $\mathbf{g}$ to into $-1\sim 1$. If adaptation prompts are initialized randomly, they may introduce noise or interfere with word tokens during the initial stages of training, thereby compromising the stability and effectiveness of the fine-tuning process.
Finally a linear projection layer calculate the output as adaptation prompts and add on decoder embedding from $\ell_{D}$th decoder layer as:
$\mathbf{E_{D}}^{\ell_{D}}=\mathbf{E_{D}}^{\ell_{D}-1}+\mathbf{Linear}(\tilde{\mathbf{S}}\mathbf{V})\in\mathbb{R}^{B\times T\times d}$ (5)
4 Experiments
In this section, we first detail datasets and implementation specifics, including MLLM feature alignment strategies and adapter training protocols (see subsection 4.1). Then, via evaluations on the OmniLabel benchmark[30], we analyze strategies for sharing vision decoder features (see subsubsection 4.2.1), investigate the impact of different adapter architectures (see subsubsection 4.2.2), and observe performance improvements across various scales of LLMs for visual detection and grounding tasks (see subsubsection 4.2.3). Finally, we conduct ablations to precisely quantify the contributions of MLLM decoder hidden states at various layers to detection performance (see subsection 4.3).
4.1 Dataset and Implementation Details
MLLMs alignment: The dataset used to train the projector follows the approach outlined in the ShareGPT-4V plan [2]. Our architecture adopts a multi-stage training paradigm, incorporating a carefully designed data composition. For InternVL2-1B, we employ a two-phase strategy:
  1. 1.
    Stage 1: This phase leverages LAION-CC-SBU (558K samples) [18] to establish cross-modal alignment through diverse visual-linguistic interactions.
  2. 2.
    Stage 2: This phase combines ShareGPT4V (102K) and SFT (665K) [2] for instruction tuning, augmented with domain-specific datasets (ChartQA [18K] [24], AI2D [12K][12], DocVQA [10K] [25]) to enhance structured visual understanding.
Stage 3 - Adapter training: We employ Grounding DINO[21] pretrain on Object365, GoldG and Cap4M as a unified detection pre-training framework across three datasets: Objects365 [31], COCO2017, and Flickr30k [38]. We implement our framework using InternVL2-1B (Qwen2-0.5B backbone) as the MLLM [4, 32, 9, 5, 6] and Open-GroundingDINO [21] as the detection codebase.
Training is conducted on 8×A100 GPUs. We pretrain (Stage 1) and fine-tune (Stage 2) the MLP layer each for 1 epoch. For pretraining, the warmup ratio, batch size, learning rate, and weight decay are set to 0, 128, $1\times{10}^{-3}$, and 0.03, respectively. The comprehensive optimization process spanned roughly 4900 steps. For fine-tuning, the warmup ratio, batch size, learning rate, and weight decay are set to 0.03, 128, $4\times{10}^{-5}$, and 0, respectively. In Stage 3,We train the Adapter with a batch size of 64 for 1 epoch. We employ cosine learning rate scheduling (base lr=$2e^{-4}$ without warmup or weight decay. The MLP layer and LLMs is trained with the adapter using a different learning rate of $4\times{10}^{-5}$.
4.2 Evaluation
We evaluated and analyzed adapters under different architectures based on the OmniLabel benchmark[30]. Additionally, we compared the performance results of the hidden state acquisition scheme in InternVL2 with GroundingDINO (abbreviated as G-DINO in the table).
4.2.1 Vision Decoder Share from Detector
The rich semantic understanding in the MLLM primarily stems from the LLM rather than the vision encoder. Therefore, we first use the architecture and weights of the vision encoder with those from GroundingDINO to provide vision features to LLMs. We shared the architecture and weights of the vision encoder for both the detector and MLLM, and discarded a series of image preprocessing schemes commonly used in MLLM. Most MLLMs employ dynamic processing that dynamically selects an optimal aspect ratio from a set of predefined ratios. This process partitions the input image into pixel-based tiles and generates a thumbnail to retain global contextual information [19, 4, 32, 9, 5, 6]. But we discarded this to eliminate the additional image encoder and avoid providing extra fine-grained features through pre-segmented images, which could confuse the LLMs’ understanding of vision tokens. We first pretrain (stage 1 in Figure 3) and then finetune (stage 2 in Figure 3) the MLLM while freezing the vision encoder and LLM during training. subsection 6.4 (See supplement) provides an MME evaluation of the alignment degree of the MLP layers in stages 1 and 2.
The vision decoder we sharing from the detector is Swin-Tiny, introduce a pixel shuffle operation with a scaling factor of 4 to downsample the spatial dimensions and expand the channel dimension. Specifically, the height $h$ and width $w$ of the vision encoder’s output are reduced by a factor of 4, and the channel dimension is increased by a factor of 16. The output is then repeated and truncated along the channel dimension to match the input dimension of the MLP projection layer (4096). The alignment of feature spaces is achieved by pretraining the MLLM’s MLP projection layer as follows:
$\hat{f}_{v}=\text{MLP}(\text{Truncate}(\text{Repeat}$ (6)
$(\text{PixelShuffle}_{4}$ $(\text{Vision Encoder}(I)))))$
where $I$ denotes the input image and $\hat{f}_{v}$ represents the visual features projected. $\text{PixelShuffle}_{4}(\cdot)$ denotes the pixel shuffle operation with a scaling factor of 4, $\text{Repeat}(\cdot)$ represents the repetition of channels, $\text{Truncate}(\cdot)$ ensures that the channel dimension matches the MLP’s input requirements.
Finally, when training the adapter (stage 3 in Figure 3), both the LLM and the MLP layers are trained simultaneously to adapt to downstream tasks. Image embeddings from Swin-Tiny (used in GroundingDINO) are fed into the LLM after alignment by a projector. The pretrained weights for Swin-Tiny, which are the same as those used in GroundingDINO and provided by Open-GroundingDINO [47], are utilized and frozen during every stages.
We measured the GFLOPs for each architecture and selected the $2_{th}$ decoder layer, removing subsequent layers. As shown in Table 1, L ED requires only 8.7% more FLOPs than GroundingDINO, with the majority of the computational cost attributed to the LLM’s decoder.
Model/Component Params (M) GFLOPS
G-DINO 172M 412G
Adapter 0.26M 1.47G
Conv Layer 2.06M 0.54G
Qwen2-0.5B $2_{th}$ 56M 34G
Total 58M 36G
Table 1: Model parameters and computational costs. It can be seen that almost all additional overhead comes from a lightweight (0.5B) LLM with only two layers, while the refinement of hidden states by the adapter is extremely efficient.
The OmniLabel benchmark [30] for our method is presented in Table 2 with an improvement of $2.33\%$ over GroungdingDINO. This improvement can be attributed to the adaptation prompts provided by the LLMs.
OmniLabel AP AP AP
descr,categ categ descr
G-DINO 21.69% 33.25% 16.09%
 Swin-T + Qwen2-0.5B 24.01% $\uparrow$2.3333.01% $\downarrow$0.2518.53% $\uparrow$2.44%
Table 2: OmniLabel evaluation of the pipeline for stage 1,2,3 and share vision decoder from detector with improvements over GroundingDINO.
4.2.2 Adapter Architecture Analysis
In MLLMs, the alignment between the vision encoder and LLMs directly determines the quality of the LLMs’ understanding of image tokens. To maximize the quality of the adaptation prompts provided by LLMs, we employ InternVL ViT and its pre-trained weights as the vision encoder for models such as InternVL2-1B, 2B, and 8B. Although the hidden states from the LLM decoder contain rich semantic understanding of the images as endowed by the LLMs, they also include a significant amount of redundant information that has already been captured by the vision encoder of the detector. In this section, we discuss various design approaches to extract adaptation prompts from the hidden states, as well as how to handle the remaining text embeddings resulting from the truncation of the hidden states.
We analyze and evaluate an optimal feature fusion strategy to design the adapter, we discussed various possible adapter design schemes:
Architecture $\mathrm{I}$ (Double Cross-Modal Fusion): As show in the Figure 6 (in the supplement), owing to the robust cross-modal feature fusion capabilities of MLLMs, the text embeddings $\mathbf{E_{T}}$ extracted from the hidden state, encompass substantial information relevant to image understanding. Furthermore, these embeddings can guide the Detector by leveraging user-provided instructions. The LLM input is formulated as a structured prompt that integrates task instructions and object captions, adhering to GroundingDINO’s annotation protocol:
<<<user>>>: [Instruction: Detect objects in this image. If present, locate them with bounding boxes.] + [Caption: {cat, dog, car,…}]‘
We extract image/text embeddings $\mathbf{E_{T}}/\mathbf{E}_{V^{D}}$ from the decoder’s hidden states using positional segmentation. The cross-attention mechanism operates through two stages:
  1. 1.
    Text-guided image fusion: Image embeddings as queries attend to text embeddings (keys/values)
  2. 2.
    Prompt-enhanced detection: Detector queries attend to fused embeddings (keys/values)
Through two cross-attention mechanisms, we integrate the Adaptation prompt, which combines images and text, back into the Image embedding of the Detector through a zero-initialized gating:
$\mathbf{A_{P}}=Gatng(CrossAtt(CrossAtt(\mathbf{E_{V}}^{L},\mathbf{E_{T}}),\mathbf{E}_{V^{D}}))$ (7)
Architecture $\mathrm{II}$ (Late Prompt Projection):We removing the cross-modal attention layers to Image embedding of the Detector. As show in the Figure 7 (See supplement), A $3\!\times\!3$ convolution (stride=2) directly process the MLLM’s fusion adaptation prompts $\mathbf{A_{P}}$to match the detector cross-modal decoder’s input space, implementing:
$\mathbf{A_{P}}=\mathrm{Conv2D}(CrossAtt(\mathbf{E_{V}}^{L},\mathbf{E_{T}}))\in\mathbb{R}^{B\times L\times d}$ (8)
Then the adaptation prompts provide to the last cross-modal decoder layer $\ell_{D}=6$, as the describle in subsection 3.4.
Architecture $\mathrm{III}$ (Early Decoder Injection): Instead of feeding prompts to the detector’s final decoder layer (Arch. II), we inject them at the first decoder layer by choosing $\ell_{D}=1$ instead of $\ell_{D}=6$ as show in the Figure 7 (See supplement). This allows gradual refinement through 6 decoder blocks.
$\mathbf{F}_{i}^{\ell_{D}}=\mathrm{Decoder}_{\ell_{D}}(\mathbf{F}_{i}^{\ell_{D}-1},\mathbf{P}),\ \ell_{D}\in\{1,...,6\}$ (9)
Architecture $\mathrm{IV}$ (Text-Free Adaptation): As shown in the Figure 4, Eliminating text embeddings $\mathbf{E_{T}}$, we directly project image embeddings $\mathbf{E}_{v}$ using the same convolutional layer as Arch. II. This tests the hypothesis that visual embeddings alone carry sufficient multimodal signals.
Figure 4: The adapter architecture IV of L ED : We ultimately used image embedding part as the key and value for cross-attention. The query is derived from the embedding of the last detector decoder. The out of attention to the input of the next decoder through a zero-initialized gating.
Comparing Arch. I and II, adaptation prompts injection into the cross-modal decoder (II) significantly improves descr-AP by 4.17% compared to I’s 0.86%, but reduces categ-AP by 2.20%. This suggests that input fusion in the cross-modal decoder preserves categorical priors, while internal fusion enhances open-vocabulary adaptation.
Arch. I and II show similar performance on the overall task, but Arch. III’s first-layer injection achieves higher categ-AP (32.07% vs. II’s 31.05%) but lower descr-AP (19.99% vs. 20.26%). We attribute this to early fusion, which enables gradual feature refinement through decoder layers, better preservation of categorical description features, and reduced overfitting to textual descriptions. Despite the rich textual understanding of MLLM (Arch. I), Arch. IV achieves superior performance (27.87% vs. 22.57% AP). This reveals two phenomena: (1) visual embeddings $\mathbf{E}_{v}$ already encode cross-modal alignment through MLLM pretraining, and (2) naive text-visual fusion may introduce semantic noise without explicit contrastive training for multimodal alignment.
Method AP AP AP
descr,categ categ descr
G-DINO 21.69% 33.25% 16.09%
I 22.57% $\uparrow$0.8833.80% $\uparrow$0.5516.95% $\uparrow$0.86%
II 24.52% $\uparrow$2.8431.05% $\downarrow$2.2020.26% $\uparrow$4.17%
III 24.63% $\uparrow$2.9532.07% $\downarrow$1.1919.99% $\uparrow$3.90%
IV 25.98% $\uparrow$4.2933.89% $\uparrow$0.6321.06% $\uparrow$4.97%
Table 3: OmniLabel evaluation for different adapter architectures. Architecture I: Full cross-attention fusion; II: Conv-based prompt projection; III: Early decoder injection; IV: Text-free adaptation.
4.2.3 Different LLMs
We employ the InternVL-1B, 2B, and 8B models to extract the hidden state from the $\ell_{LM}=$ 2nd layer of the LLM decoder to generate the adaptation prompts. InternVL-1B, 2B, and 8B models are all based on InternViT-300M as the vision encoder, and they use Qwen2-0.5B, InternLM2-1.8B, and InternLM2.5-7B as the LLMs, respectively. Table 4 demonstrates the performance comparison between our method and Grounding DINO on the OmniLabel benchmark [30].
OmniLabel AP AP AP AP
descr,categ categ descr descr-pos
G-DINO 21.69% 33.25% 16.09% 24.61%
InternVL2-1B 27.90% $\uparrow$6.2233.93% $\uparrow$0.6723.70% $\uparrow$7.6136.24% $\uparrow$11.63%
InternVL2-2B 27.87% $\uparrow$6.1933.44% $\uparrow$0.1923.90% $\uparrow$7.8136.13% $\uparrow$11.53%
InternVL2-8B 26.33% $\uparrow$4.6532.03% $\downarrow$1.2222.36% $\uparrow$6.2734.34% $\uparrow$9.73%
Table 4: OmniLabel evaluation on GroundingDINO vs. got adaptation promps InternVL-1B, InternVL-2B and InternVL-8B.
RefCOCO RefCOCO+ RefCOCOg
eval test A test B eval test A test B eval test
G-DINO 48.0% 54.8% 41.5% 48.5% 53.0% 44.1% 66.2% 66.4%
InternVL2-1B 51.3% ↑3.3% 58.6% ↑3.8% 43.4% ↑1.9% 50.8% ↑2.3% 56.0% ↑3.0% 45.3% ↑1.2% 70.2% ↑4.0% 70.5% ↑4.1%
InternVL2-8B 52.4% ↑4.4% 59.8% ↑5.0% 45.1% ↑3.6% 51.8% ↑3.3% 57.2% ↑4.2% 46.8% ↑2.7% 71.2% ↑5.0% 71.4% ↑5.0%
Table 5: Performance comparison on RefCOCO, RefCOCO+, and RefCOCOg datasets with improvements over GroundingDINO.
L
ED
achieves significant improvements in open-set object detection, with 6.22% and 6.19% AP gains for InternVL-1B and InternVL-2B, respectively. The attention mechanisms in InternVL’s decoder layers effectively map visual features to semantically rich textual embeddings, enabling better grounding of complex object descriptions. This addresses the limitations of current mainstream detectors in aligning textual descriptions with images.
In particular, the description and positioning joint task shows remarkable 11.63% and 11.53% AP improvements, demonstrating MLLMs’ capacity for spatial relationship understanding. We hypothesize that this stems from the model’s strong logical reasoning capabilities, which enhance the understanding of spatial relationships in textual descriptions, such as ”a person holding a phone.” The visual-textual co-attention mechanisms likely help disambiguate spatial configurations that challenge pure visual encoders. This phenomenon can be clearly observed in the analysis of subsection 6.7 (See supplement).
It is worth noting that category detection shows minimal improvements (0.67% AP and 0.19% gain) across experiments because closed-set category recognition relies primarily on visual prototype matching rather than semantic understanding. Thus, the Swin-Tiny encoder’s features may already saturate performance on this conventional detection task, leaving limited room for text-guided enhancement.
We evaluated our model on RefCOCO/+/g [39], and the results are shown in Table 5. Combining the evaluation results from OmniLabel, we found that: InternVL-1B shows comparable performance to InternVL-8B despite 14×14$\times$14 × more parameters. This suggests a pattern of diminishing returns in scaling LLMs for visual grounding tasks, supporting the hypothesis from [7] that small model can suffice for certain vision-language mappings.
4.3 Ablations
To quantify the different-level features contained in the hidden states obtained from various decoder layers of the MLLM, and use only the vision transformer(ViT) embeddings obtained from the MLLM’s vision encoder without the LLM ($-1_{th}$ decoder layers) to provide adaptation prompts for the detector decoder, we conducted the following ablation experiments.
Figure 5: The OmniLabel evaluation assesses adaptation prompts derived from different decoder layers in the LLM, spanning layers $-1$ to $24$. The $-1$ layer indicates that adaptation prompts are generated solely using the vision encoder in the MLLM, rather than being truncated from the hidden state.
The OmniLabel evaluation for adaptation prompts from Different Decoder Layer $\ell_{LM}$ as show in the Figure 5. From the experimental results, it is evident that the additional information covered in the adaptation prompts provided by the ViT embedding of the MLLM ($-1_{th}$ layer) alone is extremely limited and even detrimental to the performance of the categories task. However, the semantic information contained in the lower-level LLM decoder remains highly rich. In the case of InternVL2-1B, which uses a 24-layer LLM Qwen2-0.5B, utilizing just 2 layers of the LLM significantly improves the detector’s performance, peaking at the fourth layer. Notably, the adaptation prompts extracted from deeper layers of the detector do not necessarily imply higher-quality feature information. This inference fully aligns with the conclusions we derived from the attention scores in the Decoder’s hidden layers in section 3 and Figure 9 (See supplement), namely that image processing is completed in the early layers, and most of the image information has already been ”condensed” into the representations.
5 Conclusions
Our work presents a systematic framework leveraging shallow MLLM decoder layers to enhance visual grounding efficiency. Key contributions include: (1) Demonstrating that early Transformer layer hidden states retain sufficient spatial-semantic correlations for grounding; (2) Proposing a zero-initialized cross-attention adapter for effective MLLM-to-detector knowledge transfer; (3) Introducing two adaptation prompt schemes that boost descriptive target precision without degrading category accuracy. The Qwen2-0.5B with Swin-Tiny as a vision encoder scheme achieves a 2.33% OmniLabel recognition improvement (vs. GroundingDINO’s 24.01%) with only 8.7% added GFLOPs, while Qwen2-0.5B with InternVL ViT yields a 6.22% gain, still improve significant on RefCOCO/+/g. Evaluations of adapter architectures, LLM scales, and detector layers validate robustness, especially in terms of describing targets and zero-shot tasks, the improvement is significant. Our generalizable adapter design enables real-time vision-language applications while mitigating overfitting risks from manual prompt bias. This advances practical MLLM deployment in downstream tasks.
Supplementary Material
6 More Implementation Details
6.1 Adapter Architecture Detail
Architecture I (Double Cross-Modal Fusion). As show in Figure 6, this design leverages two sequential cross-attention layers to incorporate both textual and visual cues into the Detector. First, the image embeddings $\mathbf{E_{V}}_{L}$ attend to text embeddings $\mathbf{E_{T}}$, ensuring high-level semantic alignment via a text-guided image fusion. Second, the fused features further guide the Detector through a prompt-enhanced detection stage, where Detector queries attend to the fused representation. A zero-initialized gating mechanism adaptively combines these cross-modal features back into the Detector’s backbone.
Figure 6: The adapter architecture $\mathrm{I}$ of our approach.
Architecture $\mathrm{II}$, $\mathrm{III}$ (Late Prompt Projection). As show in Figure 7, we remove the second cross-attention step to streamline the flow into the Detector. A $3\times 3$ convolution directly transforms the fused adaptation prompts into the compatible feature dimension before feeding them into the final cross-modal decoder.
Figure 7: The adapter architecture $\mathrm{II}$, $\mathrm{III}$ of our approach.
6.2 Expertiment Configuration
In subsubsection 4.2.1, the experimental parameters used in stages 1, 2, and 3 for deploying our method are shown in the Table 6. In subsubsection 4.2.3, we explore how varying the parameter sizes of LLMs affect the quality of the adaptation prompt, following the experimental setup described in Table 7. In subsubsection 4.2.2, We used the Expertiment Configuration in Table 8 to test how LLM influence the quality of adaptation prompt. OmniLabel evaluation for different adapter architectures shows on Table 3. Additionally, in subsection 4.3, we evaluate the effectiveness of adaptation prompts generated by different layers of the LLM decoder, using the configuration specified in Table 9.
Expertiment config for sharing vision decoder
MLLM InternVL2-1B
$\mathbf{HD}$ from $\ell_{LM}$ 2
$\mathbf{AP}$ to $\ell_{D}$ 1
Training Dataset Object365, COCO, Flickr30k
Stage 1 $lr$ MLP: $1e^{-3}$
Stage 2 $lr$ MLP: $2e^{-5}$
Stage 3 $lr$ MLP & Adapter: $2e^{-4}$, LLM: $2e^{-5}$
adapter architecture IV
Table 6: Expertiment configuration details for vision decoder share from detector.
Expertiment config for different LLMs
MLLM InternVL2: 1B, 2B, 8B
$\mathbf{HD}$ from $\ell_{LM}$ 2
$\mathbf{AP}$ to $\ell_{D}$ 6
Training Dataset Object365, COCO, Flickr30k
adapter architecture IV
Table 7: Expertiment configuration details for different LLMs.
Expertiment config for adapter architecture
MLLM InternVL2-2B
$\mathbf{HD}$ from $\ell_{LM}$ 8
$\mathbf{AP}$ to $\ell_{D}$ 1, 6
Training Dataset COCO, Flickr30k
Table 8: Expertiment configuration details for different adapter architecture.
Expertiment config for Ablations
MLLM InternVL2-1B
$\mathbf{HD}$ from $\ell_{LM}$1,2,4,8,24
$\mathbf{AP}$ to $\ell_{D}$ 6
Training Dataset Object365, COCO, Flickr30k
adapter architecture IV
Table 9: Ablations expertiment configuration details for different decoder layer.
6.3 Dynamic Image Processing
[…]
Figure 8: Dynamic Image Processing
6.4 MLLM Alignment (Swin-T & Qwen2-0.5B)
We evaluated the alignment of Swin-T & Qwen2-0.5B on MME and compared it with the original method of InternVL2-1B (InternVL Vit & Qwen2-0.5B) as show in Table 10. To evaluate the impact of replacing the vision encoder and aligning the LLM on MLLM performance, we compared the score gaps between InternVL2-1B and Swin-T & Qwen2-0.5B on the MME benchmark. Compared to InternVL2-1B, Swin-T & Qwen2-0.5B scored 1027 on perception, a decrease of 335 (InternVL2-1B scored 1362.97), and 273.57 on cognition, a decrease of 145.72 (InternVL2-1B scored 419.29). It is worth noting that, although more advanced alignment strategies may exist, our work does not aim to achieve maximum alignment between an extremely lightweight vision encoder (Swin-Tiny) and the LLM, particularly without employing strategies such as Dynamic Tiling.
MME InternVL2-1B Swin-T & Qwen2-0.5B
total 1363 1027
existence 180.0 158.3
count 118.3 105.0
position 126.7 66.7
color 135.0 128.3
posters 110.2 75.9
celebrity 146.8 131.8
scene 148.5 120.8
landmark 132.5 83.3
artwork 140.0 80.0
OCR 125.0 77.5
total 419.29 273.57
commonsense reasoning 99.29 68.57
numerical calculation 62.50 62.50
text translation 162.50 70.00
code reasoning 95.00 72.50
Table 10: Performance comparison of InternVL2-1B, Swin-T & Qwen2-0.5B on MME Perception and Cognition tasks.
6.5 Attention Maps in Qwen2-0.5B Decoding
Figure 9: Attention maps during the decoding process of Qwen2-0.5B.
Based on the attention maps during the decoding process of Qwen2-0.5B, We draw the attention scores for different layer as fellow:
Attention $\text{Attention Score}(Q,K)=\frac{QK^{T}}{\sqrt{d_{k}}}$ (10)
Where, $Q,K$ are the Query and Key matrix in the transformer Decoder layer, $\text{softmax}(\cdot)$as show in Figure 9, we have reached an approximate conclusion [3]. We can observe that in the first layer, attention is distributed relatively smoothly across different types of tokens. In the deeper layers, starting from local attention, attention scores are aggregated onto system prompts, instructions, and output tokens, while attention to image tokens becomes quite sparse. In the deeper layers, there are strong vertical lines (in the system prompts) that dominate most of the attention scores. The presence of these strong vertical lines indicates that certain input tokens remain highly attended to throughout the decoding process. If large-scale, high-intensity attention to image tokens were also maintained in the deeper layers, it would suggest that a significant amount of visual information is still needed for inference at those stages. However, as seen in the visualization, the deeper layers primarily focus on some key text tokens, which precisely indicates that the processing of images has already been completed in the earlier layers, and most of the image information has been ”condensed” into the representations.
6.6 Dataset Introduction
Table 11 provides the attributes, sizes, and functional descriptions of all datasets used in our stages 1, 2, 3 and evaluation, to illustrate the support they offer during the process.
Table 11: Introduction of datasets used in stage 1, stage 2, and stage 3.
Dataset Introduction
Pretraining Data for Stage 1
LAION-CC-SBU 558K The LAION-CC-SBU dataset is a curated subset of the LAION, Conceptual Captions (CC), and SBU datasets. It comprises 558K image-caption pairs for the pretraining stage for feature alignment in visual instruction tuning. [18]
Instruction Tuning Data for Stage 2
ShareGPT4V ShareGPT4V dataset is curated from LAION, CC, SBU, SAM, COCO, web-landmark, web-celebrity, wikiart, etc, resulting in total 102K high-quality image-text pairs with the help of powerful GPT4-Vision [2].
SFT SFT dataset comprises approximately 665K multimodal instruction-following samples, facilitating improved alignment of visual-language models to human instructions [2].
ChartQA ChartQA is a domain-specific visual question-answering dataset containing 18K samples designed explicitly for interpreting various types of charts, including bar graphs, pie charts, and line plots. [24]
AI2D AI2D includes approximately 12K annotated diagrams paired with structured question-answer data, particularly aimed at evaluating multimodal reasoning over scientific diagrams. [12]
DocVQA DocVQA contains 10K samples that involve complex question-answering tasks over visually rich document images, emphasizing text recognition, layout analysis, and semantic comprehension. [25]
Grounding Data for Stage 3
Objects365 Objects365 is a large-scale object detection dataset featuring over 1.7 million images with dense annotations covering 365 common object categories, enhancing general object detection capabilities. [31]
COCO2017 COCO2017 provides around 118K training images annotated for object detection, and captioning tasks, widely used as a benchmark in computer vision research.
Flickr30k Flickr30k is a standard multimodal dataset with 31K images, each annotated with five descriptive captions, commonly utilized for improving image captioning and cross-modal retrieval tasks. [38]
Evaluation Data
RefCOCO/+/g The RefCOCO/+/g datasets consist of referring expression comprehension tasks, where models must identify objects in images based on natural language descriptions. REFCOCO has approximately 142K referring expressions, while RefFCOCO+ and RefCOCOg provide more challenging and generalized scenarios. [39]
Omnilabel Omnilabel is a multimodal benchmark dataset containing diverse visual-language tasks designed to comprehensively evaluate the generalization and zero-shot capabilities of visual-language models across various tasks and domains. [30]
6.7 Case Study
We compared test results on OmniLabel using InternVL2-1B for adaptation prompts against directly applying GroundingDINO on selected samples. We excluded all category-based targets and focused solely on the detection performance of descriptive targets. As shown in Figure 10, GroundingDINO struggles to understand target descriptions involving quantities, states, ages, colors, or spatial relationships. However, when the LLM provides semantic-level understanding, the detector’s ability to recognize such targets improves significantly.
Figure 10: Case Study for OmniLabel using InternVL2-1B to provide adaptation prompts versus directly using GroundingDINO.
Notably, the adaptation prompt functions as expected, guiding the detector rather than allowing the adapter to dominate the grounding task. As seen in the recognition results in Figure 10, GroundingDINO often fails to locate individuals, instead detecting entire objects, as in the case of ”The computer monitors on the desk that are turned on.” Additionally, it struggles to interpret positional relationships in text descriptions, such as ”the sandwiches that are each on the same plate with the other.” With the adaptation prompt, the detector retains an overall understanding while incorporating individual positional relationships. Furthermore, it corrects errors in spatial logic relationships.